OpenStudy (anonymous):

What is the value of the discriminant for the quadratic equation 2x2 - 3x + 1 = 0?

OpenStudy (anonymous):

Would this be 1 ?

OpenStudy (anonymous):

yes

OpenStudy (p0sitr0n):

a=2 b=-3 c= 1 b^2-4ac = 9-4*2*1] - 9-8 = 1

OpenStudy (anonymous):

Could you help me with two more questions?

OpenStudy (anonymous):

These are a bit more complex and confusing....

OpenStudy (anonymous):

You can work at most 15 hours next week. You need to earn at least $86 to cover your weekly expenses. Your job as a waiter pays $7.25 per hour, and your job as a dog walker pays $6.75 per hour. Identify the system of linear inequalities that models the situation where x is the number of hours worked as a waiter, and y is the number of hours worked as a dog walker. x + y >= 15 7.25x + 6.75y <= 86 x + y <= 15 7.25x + 6.75y >= 86 x + y <= 15 6.75x + 7.25y >= 86 x + y >= 15 6.75x + 7.25y <= 86

OpenStudy (anonymous):

You earn 7.25 as a waiter, and work x hours as a waiter, so you earn 7.25x as a waiter You earn 6.75 as a walker, and work y hours as a walker, so you earn 6.75x as a walker Total earnings = 7.25x + 6.75y Your earnings must total at least $86 So 7.25x + 6.75y ≥ 86 Given you can only work upto 15 hours per week, you know that x + y ≤ 15 So your answers are: 7.25x + 6.75y ≥ 86 and x + y ≤ 15

OpenStudy (p0sitr0n):

x: waiter y: dog x+y=<15 (less than 15, but can be 15) 7.25*x+6.75*y >= 86

OpenStudy (anonymous):

Ok, that makes sense... one more please

OpenStudy (anonymous):

With a tail wind, a plane traveled 800 mi in 5 hours. With a head wind, the plane traveled the same distance in 8 hours. Which system of equations could be used to solve this problem? (s + w)5 = 800 and (s - w)8 = 800 (s - w)5 = 800 and (s + w )8 = 800 5s + 8w = 800 and 5s - 8w = 800

OpenStudy (anonymous):

well well working for free thats not acceptable ...

OpenStudy (anonymous):

I'm pretty sure it's (s + w)5 = 800 and (s - w)8 = 800 I just wanted someone to double check but nvm I guess

OpenStudy (anonymous):

Thanks anyways!